What is the Molality of a solution containing 7.78g of Urea [(NH2)2CO2] in 203g of water? 


1
Expert's answer
2021-06-02T06:31:58-0400

"Molarity= moles\/litre"

Moles = 7.78/60 =0.1297moles

Molarity= (0.1297× 0.203)/1000

= 2.63×10-3M